# Asymmetry Analysis of Bilateral Shapes

### v1

@@ 5 Discussion
We have presented our work for pre-registered data, but it can be extended when the landmark data is not pre-registered as pointed out in Section 4. Further, these revised measures of asymmetry can be applied and analyzed in a straightforward way. Further, we have used only size-and-shape framework but it can be extended easily to include for example, similarity shape.
We have used absolute values ($a_{j}$) on coordinate elementary features as in equation (2.5) which is of main interest if one is looking for the extent of asymmetry as in our case for the smile data. However, there maybe some other applications where raw $d_{j}$ may be important than $a_{j}$; our work easily applies to that situation.
In this paper we have concentrated on smile example with three fixed frames. Our future statistical work will be extended to dynamic shape analysis which will allow to study trajectories of the smile. Some initial work for a single group of a smile data has been introduced in Mardia et al. (2018) and Bookstein et al. (2024). We have concentrated on the cleft lip subjects but there are other medical applications such as in facial reconstruction related to trauma surgery where the lips form significant area for surgery.
We have used univariate methods but we can use Hotelling’s $T^{2}$ in place of t-statistics, for example, though it requires some modifications. An one-sided alternative seems natural according to equation (3.1) and in future we would consider one-sided multivariate tests. Nevertheless, it requires more care to set up than in the one-dimensional case. Further, the scenario of sample size is less than the dimension of the data vector has also to be taken into account.
Acknowledgements
The authors would like to thank the School of Mathematics, University of Leeds for supporting the smile project. Kanti Mardia acknowledges the Leverhulme Trust for the Emeritus Fellowship.
Appendix A Estimation of midplane using Procrustes analysis
Let $\bm{n}\in\mathbb{R}^{M}$ denote the unit normal vector of an arbitrary plane $\mathcal{P}^{\prime}$. Let $X_{n}\in\mathbb{R}^{K\times M}$, $n=1,\dots,N$ denote the observed configurations. The reflection of $X_{n}$ about $\mathcal{P}^{\prime}$, $X^{\text{(refl)}}_{n}$, is given as
$X^{\text{(refl)}}_{n}=X_{n}H,$
where $H=I_{M}-2\bm{n}\bm{n}^{T}$ is the Householder matrix and $I_{M}$ is the $M\times M$ identity matrix. Then the estimation process used in basis registration is given as the following:
  1. 1.
    Form an augmented dataset: $\{X_{n},X^{\text{(refl)}}_{n}\}_{n=1}^{N}$.
  2. 2.
    Apply GPA on this augmented dataset. Denote the fitted configurations as $X_{n}^{(\text{GPA})}$ for $n=1,\dots,N$.
  3. 3.
    Compute the Procrustes mean shape. This mean shape will be bilaterally symmetric with midplane $\mathcal{P}^{\prime}$.
  4. 4.
    Rotate the mean shape outside and within $\mathcal{P}^{\prime}$ until the new midplane and the basis are meaningful.
  5. 5.
    Carry out ordinary Procrustes analysis between each $X_{n}^{(\text{GPA})}$ and the rotated mean shape.
When axis registration is considered, we only rotate the sample mean shape outside midplane. Further, the last step, step 5, is ignored. Further details are given for this procedure in Mardia et al. (2000).

### v2

@@ 5.4 Summary of the Analyses
6 Analyses of the Orthognathic Surgery Smile Data
6.1 Details
The orthognathic surgery smile data contains $N=22$ subjects who have undergone orthognathic surgery and their pre- and post-surgery data have both been collected. Figure 7 shows the face before (in the left) and after (in the right) the orthognathic surgery, where the view is from the right side of the face. It can be seen in the left of Figure 7, the jaw is somewhat protruded in comparison to the figure on right.
Figure 7: View of the face from right before (left) and after (right) the orthognathic surgery.
The axes used for the Smile Data 2 are exactly the same as the Smile Data 1 and the details are given in Section 2.1. The 16 landmarks on the nasal region are shown in left of Figure 8, while the right of Figure 8 shows the 10 landmarks on the lip. The landmark indices are summarized in Table 7. The scale of the Smile Data 2 is in mm. The same frames as the Smile Data 1 are used.
Table 7: Indices of landmark pairs and solos in Figure 8 on the lip periphery and nasolabial region for the Smile Data 2.
Landmark notation Indices in Figure 8
pair $(k_{L},k_{R})$ (1,7), (2,6), (3, 5), (8,10), (11,12), (15,16)
$k_{S}$ 4, 9, 13, 14
There are two sets of landmarks we will focus on:
  1. (a)
    The 7 landmarks on the upper lip (landmarks 1 to 7 in Figure 8) and the 6 landmarks on the nasal region (landmarks 11 to 16 in the left of Figure 8). There are $K_{P}=5$ paired landmarks and $K_{S}=3$ solo landmarks, hence, $J=18$.
  2. (b)
    The 10 landmarks on the lip (landmarks 1 to 10 in Figure 8). There are $K_{P}=4$ paired landmarks and $K_{S}=2$ solo landmarks, hence, $J=14$.
Figure 8: The left figure shows the 16 landmarks on the nasolabial region for a typical subject at the first frame using the pre-surgery data. The right figure shows the 10 landmarks on the lip periphery. The two figures are based on the Smile Data 2.
These two sets of landmarks correspond to two aims:
  1. (a)
    Assess the affects of surgery, by comparing pre- and post-surgery data via paired tests (item (a) above).
  2. (b)
    Compare the subjects from Smile Data 2 to the control subjects from the Smile Data 1 where we use the same landmarks on the lip for both data (item (b) above).
Some previous studies have been carried out, such as in Al-Hiyali et al. (2015), Xue et al. (2020), Xue et al. (2023) and Quast et al. (2023). These studies have focused on different facial expressions and the authors have taken average over measurements resulted from these expressions. However, by doing this, they would lose a lot of information. Our work is different from these previous studies as we attempt to construct asymmetry measurements over the lip and nasal region at different frames during the smile.
The rest of this section will be divided into two parts corresponding to item (a) (effect of surgery, see Section 6.2) and item (b) (deviation from normal subjects, see Section 6.3) above.
6.2 Test Results for Smile Data 2: Effect of Surgery
The pre-surgery data is compared to the post-surgery data by using one-sided paired tests on the asymmetry measurements. Let $g=1,2$ denote the pre-surgery data and post-surgery data respectively. The absolute elementary feature vectors $\bm{a}_{n}^{(1)}$ and $\bm{a}_{n}^{(2)}$ are computed using the pre- and post-surgery data separately, for $n=1,\dots,22$.
6.2.1 Application of the Combine-then-compare Approach
The 13 landmarks on the upper lip and nasal region are used. Paired t-tests have been performed on composite asymmetry scores defined in equation (3.9) whereas Wilcoxon signed-rank tests have been performed on scores defined in (3.6), (3.7) respectively.
Since the expert suspects that the asymmetries are increased after the surgery, so the one-sided alternative hypothesis becomes the distribution relates to pre-surgery data is stochastically smaller than that of post-surgery data. The Group 1 and Group 2 in (3.1) are changed to pre- and post-surgery data respectively.
The test results on $\phi^{*}_{L_{1}}$ given in equation (3.9) are shown in Table 8. $H_{0}$ is rejected in middle and last frames. In other words, the asymmetries are increased after surgery when inspect at these two frames. Further, the $p$-value is the most extreme at the last frame, which indicates that the two groups are the most different at the last frame.
Table 8: Mean, variance, t-values and $p$-values using $\phi^{*}_{L_{1}}(\bm{a})$ based on the Smile Data 2 (equation (3.9)).
Pre Post One-sided paired test
$\phi^{*}_{L_{1}}(\bm{a})$ mean sd mean sd $t$-values $p$-values
First 15.69 5.64 18.47 8.49 -1.33 0.098
Middle 17.53 5.92 22.25 10.89 -2.36 0.014(*)
Last 17.72 5.45 22.77 9.52 -3.29 0.002(**)
Table 9 displays the results of Wilcoxon signed-rank tests on $\phi_{L_{1}}(\bm{a})$ given by (3.6) and $\phi_{L_{2}}(\bm{a})$ given by (3.7). We reach the same outcomes as the paired t-tests: $H_{0}$ is rejected at the middle and last frames and the two groups are the most different at the last frame.
Table 9: $p$-values from one-sided Wilcoxon signed-rank tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7) based on the Smile Data 2.
Composite asymmetry score First frame Middle frame Last frame
$\phi_{L_{1}}(\bm{a})$ 0.13 0.016(*) 0.002(**)
$\phi_{L_{2}}(\bm{a})$ 0.088 0.007(**) 0.001(**)
6.2.2 Application of the Compare-then-combine Approach
Feature selection The procedures given in Section 3.2 are performed using pre- and post-surgery data, in order to select which landmarks contribute the most to the increment in asymmetries after the surgery. 10000 bootstrap iterations are performed. The $25\%$, $20\%$, $15\%$ and $10\%$ critical values of $V$ are determined at the three frames separately. The corner pair, landmarks (1, 7) are important at the middle frame with respect to $25\%$, $20\%$ and $15\%$ critical values of $V$. The significance levels used here are larger than we used for Smile Data 1 in Section 5, which implies the features selected here are less important than the features in Smile Data 1.
Meta-analysis Fisher’s method and Pearson’s method given in (4.1) and (4.2) are performed with Wilcoxon signed-rank tests as individual tests using $\bm{a}_{n}^{(1)}$ and $\bm{a}_{n}^{(2)}$, for $n=1,\dots,22$. $P_{L}$ defined in equation (4.1) is computed and compare with the tail of $\chi^{2}_{36}$ distribution (as $J=18$). As for $Q$, permutation tests are used to estimate the overall $p$-values. Table 10 shows the values of $P_{L}$, $P_{R}$, $Q$ and $p$-values for Fisher’s and Pearson’s method at the three frames.
Table 10: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames. Pre- and post-surgery data from the Smile Data 2 is used here.
$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
First frame 45.52 20.04 45.52 0.13 0.45
Middle frame 63.39 10.09 63.39 0.003(**) 0.085
Last frame 64.20 11.53 64.20 0.003(**) 0.041(*)
For the Fisher’s method, we obtain the same results as in the combine-then-compare approach: $H_{0}$ is rejected at middle and last frames. However, the $p$-values obtained in Fisher’s method are more extreme at the middle frame than those from combine-then-compare approach. For the Pearson’s method, we fail to reject $H_{0}$ at the first and middle frames. This indicates that the Pearson’s method may not be sensitive to changes in asymmetries before and after the surgery.
6.3 Test Results for Smile Data 1 and Smile Data 2: Deviation from Control Subjects
The Smile Data 2 is compared with the control subjects from the Smile Data 1. In Section 5, we only use 12 subjects which form a subset of the control group. The expert believes that more data should be used, thus, all 109 control subjects are used in this section. To make the comparisons available, a subset of 10 landmarks on the lip (Figure 8) are chosen from the 24 landmarks (see Figure 3), so that the Smile Data 2 and the Smile Data 1 have the same landmarks at the lip. Hence, the absolute elementary feature vectors $\bm{a}_{n}$ for $n=1,\dots,12$ are recomputed on the subjects from Smile Data 1 using the new landmarks. Pre- and post-surgery data are used separately. Hence, we have three groups in total, but we stick with two-sample tests. We scale the data so that both datasets have the same scale in cm.
6.3.1 Application of the Combine-then-compare Approach
The one-sided two-sample t-tests are performed on $\phi^{*}_{L_{1}}$ defined in (3.9) and the outcomes are shown in Table 11. Table 12 shows the means and variances of $\phi^{*}_{L_{1}}$ for the three groups of data. $H_{0}$ is rejected in almost all cases, except for the pre-surgery data at middle frame. This implies that in most of the cases, control subjects from the Smile Data 1 are more symmetric than subjects from the Smile Data 2 .
Table 11: $p$-values from one-sided two-sample Welch’s $t$-test using $\phi_{L_{1}}^{*}$ defined in (3.9) to compare between Smile Data 2 subjects with normal subjects from the Smile Data 1.
$\phi_{L_{1}}^{*}$ First frame Middle frame Last frame
Pre-surgery 0.006(**) 0.073 0.031(*)
Post-surgery 0.025(*) 0.013(*) 0.010(*)
Table 12: Mean, standard deviation, t-values of $\phi^{*}_{L_{1}}(\bm{a})$ (equation (3.9)) for pre-, post-surgery subjects from Smile Data 2 and 12 control subjects from Smile Data 1.
Pre Post Control
$\phi^{*}_{L_{1}}(\bm{a})$ mean sd mean sd mean sd
First 0.95 0.35 0.87 0.29 0.74 0.22
Middle 1.17 0.41 1.30 0.52 1.03 0.35
Last 1.22 0.34 1.32 0.46 1.07 0.33
One-sided two-sample Mann-Whitney U tests have been performed using $\phi_{L_{1}}$ and $\phi_{L_{2}}$ defined in (3.6) and (3.7) respectively. The test results are shown in Table 13. $H_{0}$ is rejected in almost all cases except for pre-surgery data at the middle frame and post-surgery data at the first frame.
Table 13: $p$-values obtained from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}$ in (equation (3.6)) and $\phi_{L_{2}}$ in (3.7) to compare between Smile Data 2 subjects with control subjects from the Smile Data 1.
Pre Post
$\phi_{L_{1}}$ $\phi_{L_{2}}$ $\phi_{L_{1}}$ $\phi_{L_{2}}$
First 0.0044(**) 0.0050(**) 0.059 0.070
Middle 0.062 0.060 0.0081(**) 0.014(*)
Last 0.018(*) 0.040(*) 0.0030(**) 0.0037(**)
However, the test results violate the expert’s observations: both pre- and post-surgery data are asymmetric at the first frame and the pre-surgery data is more asymmetric at this frame. We then compute the $\phi_{L_{1}}$ and $\phi_{L_{2}}$ with different weights for the solos, where the weights for the solos are the doubled unit lengths ($w_{j}=2$ in (3.6) and (3.7)) while we still keep the weights for landmark pairs to be unit length ($w_{j}=1$ in (3.6) and (3.7)). The test results from Mann-Whitney U tests are displayed in Table 14. $H_{0}$ is rejected for both groups at the first frame. Further, the $p$-values for pre-surgery group is more extreme, which indicates that the pre-surgery data are more asymmetric. This findings match the expert’s opinion. Note that changing the weights in this way do not affect the results for compare-then-combine approach.
Table 14: $p$-values obtained from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}$ in (equation (3.6)) and $\phi_{L_{2}}$ in (3.7), where $w_{j}=1$ for pairs and $w_{j}=2$ for solos, to compare between Smile Data 2 subjects with control subjects from the Smile Data 1.
Pre Post
$\phi_{L_{1}}$ $\phi_{L_{2}}$ $\phi_{L_{1}}$ $\phi_{L_{2}}$
First 0.0047(**) 0.0035(**) 0.026(*) 0.018(*)
Middle 0.073 0.075 0.0055(**) 0.0073(**)
Last 0.015(*) 0.042(*) 0.0018(**) 0.0026(**)
The patterns of the $p$-values shown in Table 11, 13 and 14 are the same: the pre-surgery data has smaller $p$-values at the first frame, whereas the post-surgery data has more extreme $p$-values at the middle and last frames.
6.3.2 Application of the Compare-then-combine Approach
Feature selection The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information for the pre- and post-surgery data when compare with the normal subjects. $v_{j}$ is selected to be the two-sample t-statistics (equation (3.15)). $V$ defined in (3.16) is computed and its critical value at $5\%$ significance level is determined via bootstrap. The total number of bootstrap iterations used is 10000.
The central pair, landmarks (3, 5), possesses t-values which exceed the critical values of $V$ in most of the cases for both pre- and post-surgery data. The solo landmark 9 also corresponds to t-values exceed the critical values of $V$ in some cases. Most of the extreme p-values obtained from Mann-Whitney U tests pertain to the central pair 3, 5 and two solos, 4 and 9, where $\phi_{L_{1}}$ and $\phi_{L_{2}}$ are used. Landmark pair (3, 5) is the same as landmark pair (6, 8) and solo landmarks 4 and 9 are the same as landmarks 7 and 19 in Figure 4, so we recover the ‘Y’-shape as in Section 5.
Meta-analysis The Fisher’s and Pearson’s methods given in (4.1) and (4.2) are executed to compare the subjects from Smile Data 2 with control subjects from the Smile Data 1. $P_{R}$ defined in equation (4.1) is computed at the three frames separately and compared with the tail of $\chi^{2}_{28}$ (as $J=14$).
Pre-surgery data Table 15 shows the results at the three frames which are obtained by comparing pre-surgery data with control subjects. Comparing with Table 11 and 14, $H_{0}$ is also rejected at the middle frame for the Fisher’s method and the $p$-values are more extreme at the first and last frames. The reason could be that the means of some asymmetry features are different for the two groups. However, these features have small magnitudes, so they do not contribute significantly towards the values of composite scores. Consequently, they do not affect the test results on the composite scores, but impact the outcomes of Fisher’s method. Nevertheless, for the Pearson’s method, the results are the same as in Table 11 and 14 and the magnitudes of $p$-values are similar as well.
Table 15: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames. Pre-surgery data from the Smile Data 2 is compared with the control subjects from the Smile Data 1 here.
$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
First frame 9.54 67.13 67.13 $4.63\times 10^{-5}$(***) 0.003(**)
Middle frame 22.25 46.44 46.44 0.016(*) 0.11
Last frame 17.16 51.90 51.90 0.0039(**) 0.038(*)
Post-surgery data Table 16 shows the results obtained by comparing post-surgery data and control subjects. We obtain the same test results as in Table 11 and 14. The magnitudes of $p$-values for Pearson’s method are similar with the $p$-values obtained for $\phi_{L_{1}}$ and $\phi_{L_{2}}$ at the middle and last frames, whereas the $p$-value is decreased at the first frame for Pearson’s method. For Fisher’s method, the $p$-values are more extreme.
Table 16: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames. Post-surgery data from the Smile Data 2 is compared with the control subjects from the Smile Data 1 here.
$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
First frame 14.26 67.61 67.61 $3.98\times 10^{-5}$(***) 0.003(**)
Middle frame 10.83 70.64 70.64 $1.51\times 10^{-5}$(***) 0.004(**)
Last frame 13.64 63.86 63.86 0.00013(***) 0.009(**)
6.4 Summary of the Analyses
We end the section by summarizing our findings on the orthognathic surgery data:
  • (a)
    The asymmetries are increased after the surgery at middle and last frames.
  • (b)
    The subjects contained in the Smile Data 2 are more asymmetric than the control subjects from the Smile Data 1, except for the pre-surgery data at the middle frame.
  • (c)
    The central pair and solo landmarks, i.e. the ‘Y’-shape in Figure 4, carry important asymmetry information when we compare the subjects in Smile Data 2 with control subjects from Smile Data 1.
@@ 7 Discussion
We have presented our asymmetry measures for the registered Smile Data. But these asymmetry measures are available to any registered dataset. Further, we have used only a size-and-shape framework but the composite measures can be extended easily to include for example, similarity shape, since once we have registered the shapes, the asymmetry measures are applicable after allowing for scaling.
We have used absolute values ($a_{j}$) on coordinate elementary features as in equation (2.5) which is of main interest if one is looking for the extent of asymmetry as in our cases for the illustrative smile examples. However, there maybe some other applications where raw $d_{j}$ may be important than $a_{j}$, such as when analyzing the trajectories of the landmarks; our work easily applies to that situation.
In this paper we have concentrated on the illustrative smile examples with three fixed frames and our future statistical work will be extended to dynamic shape analysis which will allow us to study trajectories of the smile.
In our case, the number of landmarks $K$ is small but when $K$ is large we can study the distribution of asymmetry measure $\phi_{L_{2}}$ under isotropic normality and various tests can be then performed. The work leads to understanding for the sum of weighted $\chi^{2}$-distributions. Mardia and Wu (2025) studies this approach.
We have used mostly univariate methods except for the Hotelling’s $T^{2}$ in Section 5, though in our case, the sample size is less than the dimension of the elementary feature vector so it is not appropriate as Hotelling’s $T^{2}$ is singular. Even in the case when the sample size is large enough, the one-sided alternative seems natural according to equation (3.1). Such cases are also part of our future investigation.
We note that the final aim of our work is to provide in future a smile index based on these asymmetry measures to be used as a standard index in clinics.
Data Availability Statement
The codes and data for this paper are available at the following website: https://github.com/XW-2025-hub/JRSS-C-Supplementary-Materials.
Acknowledgments
The authors would like to thank the School of Mathematics, and the School of Dentistry University of Leeds for supporting the smile project; we are also grateful to Fred Bookstein for collaborating with this project. The authors would also like to thank Sofya Titarenko for her helpful comments. Kanti Mardia acknowledges the Leverhulme Trust for the Emeritus Fellowship.